express log2 1/6=-3 in exponential form
Answer by stanbon(75887)

express log2 1/8 = -3 in exponential form
------
log = exponent = -3
base = 2
result = (1/8)
------------------
2^(-3) = 1/8
================
Cheers,
Stan H.